OpenStudy (anonymous):

HELP ! WILL MEDAL>> Given the general form of the equation of a circle x^2=10x+y^2-4y+20=0 *Find the corresponding form * Identify the center and radius of the circle *Find the x,y intercepts *Graph the circle I am in need of help with this asap .

OpenStudy (irishboy123):

you gotta complete two squares here.... does that mean anything to you?!?!

OpenStudy (anonymous):

No this is why I am asking for help.

OpenStudy (anonymous):

hi

OpenStudy (irishboy123):

is it this? \( x^2-10x+y^2-4y+20\) there are 2 ='s in your OP

OpenStudy (irishboy123):

mm \(x^2-10x+y^2-4y+20 \color{red}{= 0}\)

OpenStudy (triciaal):

equation of a circle (x-h)^2 + (y-k)^2 = r^2 where (h,k) is the center and r is the radius

OpenStudy (triciaal):

can you regroup the original and complete-the-square for the 2 sections?

OpenStudy (triciaal):

Is that a yes and I am working on it or is that a no I am not here so I am not responding?

OpenStudy (anonymous):

I am working on it .Math is not my best subject I'm sorry I am terrible at it :/

OpenStudy (mathmale):

Why shoot yourself in the foot? "I am working on it .Math is not my best subject I'm sorry I am terrible at it :/" sounds SOOO negative. Why not think of math as just another challenge and give y ourself credit for being able, with practice, to be able to understand it.

OpenStudy (mathmale):

In short, we have to take steps toc onvert x^2=10x+y^2-4y+20=0 into an equation of the form\[\frac{ (x-h)^2 }{ a^2 }+\frac{ (y-k)^2 }{ b^2}=1\]

OpenStudy (mathmale):

and it happens that "completing the swuare" is one method, perhaps the best one, for accomplishing that. Please, share info with the rest of us regarding what you do and what you don't understand, so that helpers can start at a point that is right for you.

OpenStudy (mathmale):

@celli2013: Why not start over? Post the following question: "I need to re-write x^2=10x+y^2-4y+20=0 in the standard form for the equation of a circle. Would someone please guide me through the necesssary "completing the square?"